### v1

# High-precision measurement of the atomic mass of 84Sr and implications to isotope shift studies

Abstract
The absolute mass of 84Sr was determined using the phase-imaging ion-cyclotron-resonance technique with the JYFLTRAP double Penning trap mass spectrometer. A more precise value for the mass of 84Sr is essential for providing potential indications of physics beyond the Standard Model through high-precision isotope shift measurements of Sr atomic transition frequencies. The mass excess of 84Sr was refined to be -80649.229(37) keV/c2 from high-precision cyclotron-frequency-ratio measurements with a relative precision of 4.810-10. The obtained mass-excess value is in agreement with the adopted value in the Atomic Mass Evaluation 2020, but is 30 times more precise. With this new value, we confirm the previously observed nonlinearity in the study of the isotope shift of strontium. Moreover, the double-beta () decay value of 84Sr was directly determined to be 1790.115(37) keV, and the precision was improved by a factor of 30.
I Introduction
Isotope shifts in atomic transition frequencies arise from differences in neutron numbers among isotopes sharing the same atomic number. The isotopic shifts in the frequency of an atomic transition demonstrate an approximately linear correlation with the isotope shift observed in a second transition. These shifts reveal contributions from field and mass shifts king1984isotope, originating from the differing nuclear masses of isotopes and variations in their nuclear charge distribution. Isotopic shifts can be systematically studied using a King plot analysis. In this analysis, the isotope shifts in two sets of electronic transitions within the same isotopes are correlated. To perform the King plot analysis, one measures the energies of two transitions for three or more isotopes of a specific element. The King plot is expected to exhibit linearity, with the experimentally determined slope serving as a reliable benchmark for theoretical predictions flam18. Deviations from linearity are crucial for refining atomic structure calculations dammalapati2009; shi2017; naze2015. Recent theoretical proposals suggest that the linearity in King plots could be leveraged to impose constraints on higher-order effects on isotope shifts or on physics beyond the Standard Model (SM) frugiuele2017; berengut2018. Motivated by this, a recent surge in efforts miyake2019; manovitz2019; knollmann2019; counts2020; solaro2020; NESTERENKO2020-2; Counts20; rehbehn2021; muller21; hur2022; figueroa2022; ono2022 has significantly advanced the precision of isotope shift measurements. Essential to placing constraints on proposed electron-neutron interactions and other novel physics through King’s linearity is a two-fold requirement. First, experimental data in the form of precision optical spectroscopy and nuclear mass measurements are needed to empirically limit the potential size of the nonlinearity. Second, if a nonlinearity is observed, precise atomic and nuclear theory is necessary to calculate beyond-first-order SM sources of nonlinearity. A natural inquiry emerges regarding the potential for the King plot to maintain its linearity at an enhanced level of experimental precision. Recent experiments conducted with strontium and ytterbium ions miyake2019; Counts20 have provided initial indications that this linearity is, in fact, disrupted at a magnitude of several standard deviations. Strontium presents favorable properties for studying isotope shifts, boasting an abundance of stable isotopes and very narrow optical transitions stellmer2014. Earlier theoretical work has also proposed the measurement of strontium isotope shifts as a promising probe for new physics frugiuele2017; berengut2018. Strontium features four stable isotopes: three bosons (84,86,88Sr) and one fermion (87Sr). Precise and accurate determination of the atomic mass of the stable Sr isotopes is crucial for probing potential causes for such nonlinearities.
Atomic masses of 86-88Sr have been measured at the FSU Penning trap rana2012 with high precision (uncertainty of 6 eV/c2) as adopted in the most recent Atomic Mass Evaluation 2020 (AME2020) Wang2021; Kondev2021. The atomic mass uncertainty for 84Sr, the least abundant naturally occurring isotope of strontium, is however notably higher at 1.2 keV/c2. For investigating fundamental physics and exploring phenomena beyond the SM via high-precision King-plot tests with strontium, it is critical to measure the atomic mass of 84Sr directly with high precision. In this article, we report on the most precise absolute mass value of 84Sr to date, achieved via high-precision cyclotron-frequency-ratio measurements with the JYFLTRAP double Penning trap mass spectrometer (PTMS). We employ this refined mass value in an updated King plot analysis for two transitions of strontium.
II Experimental method
The experiment was performed at the Ion Guide Isotope Separator On-Line facility (IGISOL) using the JYFLTRAP double PTMS Eronen2012a; Eronen2016, at the University of Jyväskylä Moore2013. Stable 84Sr+ ions were generated using an offline glow-discharge ion source. For a precise mass measurement of 84Sr, reference ions of 84Kr+ with a well-known mass value (uncertainty of 4 eV/c2 Wang2021) were concurrently produced from the same ion source. As illustrated in Fig. 1(a), the gas cell within the glow-discharge ion source comprises two sharp electrodes, with one composed of naturally abundant strontium. By introducing a gas containing natural krypton, it facilitates the simultaneous generation of stable ions for both strontium and krypton.
The generated ions of 84Sr+ and 84Kr+ were extracted using helium gas flow and electric fields facilitated by a sextupole ion guide (SPIG) Karvonen2008. Following acceleration over an electric potential of 30 kV, the ions with mass number of = 84 were mass-separated using a 55∘ dipole magnet with a mass resolving power of 500. Post isobaric separation, the ions were directed to a radiofrequency quadrupole cooler-buncher (RFQ) Nieminen2001, where they were accumulated, cooled, and bunched.
The ion bunches from the RFQ were injected to the JYFLTRAP double PTMS, consisting of two cylindrical Penning traps equipped with a 7-T superconducting solenoid. The first trap, functioning as a purification trap, is filled with buffer gas and is employed for isobaric purification using the sideband buffer gas cooling technique Savard1991. This technique alone achieves mass purification with a resolving power of approximately by selectively converting ion motion from magnetron to reduced cyclotron motion.
In the purification trap, all cooled and centered ions (84Sr+ and 84Kr+) were initially excited to a large orbit of revolution by applying a dipole excitation at the magnetron motion frequency for approximately 11 ms. Subsequently, a quadrupole excitation was executed to center the ions of interest (only 84Sr+ or 84Kr+) through collisions with the buffer gas for about 100 ms.
In the precision trap, the phase-imaging ion-cyclotron-resonance (PI-ICR) method Nesterenko2018; nesterenko2021 was employed to measure the actual cyclotron frequency, , where is the magnetic field strength, is the charge state, and is the mass of the stored ion. The scheme of the PI-ICR technique nesterenko2021; Nesterenko2018; Eliseev2014; Eliseev2013 at JYFLTRAP relies on direct measurements of the cyclotron motion and magnetron motion simultaneously by projecting their radial ion motion onto a position-sensitive MCP detector.
To determine the phases of the radial motions, the center has to be determined for the ion spots on the detector. This is done by storing the ions for a few milliseconds without exciting their cyclotron motion, and the ions are directly extracted from the trap and projected onto the MCP detector.
Two patterns, as detailed in nesterenko2021; Nesterenko2018, are utilized to measure the magnetron or cyclotron motion phases, respectively. The angle between two phase images of the projected radial motions with respect to the center spot is , where and are the polar angles of the cyclotron and magnetron motion phases. The cyclotron frequency is derived from:
where is the full number of revolutions of the measured ions during the phase accumulation time . A few different accumulation times for 84Sr+ and 84Kr+ were used to confirm the unambiguity of the cyclotron frequency. A fixed accumulation time of 400 ms was employed for the actual measurements to determine the final . A measurement with "cyclotron" and "magnetron" phase spots collected with respect to the center spot is shown in Fig.1(b).
The atomic mass was derived from the measured cyclotron frequency ratio ( = ) measurements of singly charged ions of the decay pair 84Sr-84Kr:
Here, and represent the respective masses of the parent and daughter atoms, and denotes the charge state for singly charged ions (). and correspond to the mass of an electron and the speed of light in vacuum. The electron binding energies, and , are 5.69486745(12) eV and 13.9996055(20) eV, respectively, as obtained from NIST_ASD.
The value for the double-beta decay of 84Sr can be determined from the mass difference: .
| ME | |||
| This Work | 1.000 022 902 36(48) | 1790.115(37) | -80649.229(37) |
| AME2020 | 1789.8(12) | -80649.6(12) |
III Results and discussion
III.1 Mass and -value determination
A full scanning measurement (one cycle) of the magnetron phase, cyclotron phase, and center spot in sequence was completed in less than 3 minutes for each ion species of 84Kr+ and 84Sr+. In the analysis, the position of each spot was fit using the maximum likelihood method. A few cycles were summed to ensure reasonable counts for fitting before determining the position of each spot. The phase angles were calculated accordingly based on the determined positions of the phases to deduce the cyclotron frequencies of each ion species. The cyclotron frequency of the daughter ion 84Kr+ was used as a reference and was linearly interpolated to the time of the measurement of the parent 84Sr+ (ion of interest) to deduce the cyclotron frequency ratio . Bunches with less than five detected ions per bunch were considered in the data analysis to reduce a possible cyclotron frequency shift due to ion-ion interactions Kellerbauer2003; Roux2013. Up to 5 detected ions per bunch were taken into acount for the analysis, and no count-rate related frequency shifts were observed in the analysis. The temporal fluctuation of the magnetic field for JYFLTRAP was measured to be /min nesterenko2021, where is the time interval between two consecutive reference measurements. The contribution of temporal fluctuations of the magnetic field to the final frequency ratio uncertainty was less than since the parent-daughter measurements were interleaved with minutes. To minimize the systematic uncertainty arising from the conversion of cyclotron motion to magnetron motion and potential distortion of the ion-motion projection onto the detector, the positions of the magnetron-motion and cyclotron-motion phase spots were deliberately chosen. The angle between them was kept to a few degrees or less, effectively reducing this uncertainty to a level well below 10-10 Eliseev2014. Moreover, the commencement of the initial dipolar excitation with frequency was systematically scanned across one magnetron period (6 points), and the extraction was scanned over one cyclotron period (6 points) to mitigate any lingering effects of residual magnetron and cyclotron motion that might influence the distinct spots. The measurements were conducted in eight separate time slots, each lasting around 4 hours. For each slot of measurement, a weighted mean ratio was calculated, and the maximum of internal and external errors Birge1932 was selected. The final ratio was then obtained as a weighted mean ratio of all sets, taking into account the maximum of internal and external errors. The systematic uncertainties and experimental statistical uncertainty were added in quadrature to the final result. The final frequency ratio with its uncertainty, as well as the corresponding and mass-excess values, are 1.000 022 902 36(48), 1790.115(37) keV, and -80649.229(37) keV/c2 respectively.
In Fig.2, the analysis results, which include all data sets, are compared to literature values. A comparison of these results to the literature values are also tabulated in Table. 1.
The mass excess (-80649.229(37) keV/c2) and (1789.8(12) keV) from this work are both a factor of 30 more precise than those derived from the evaluated masses in AME2020 Huang2021; Wang2021, but both agree well with the values documented in AME2020. The mass-excess value in AME2020 is derived primarily from two PTMS experiments Keller07; Emma11 with an influence of 88.8%. A slight contribution of 6.8% is from endpoint energy measurements of 84Rb()84Sr BOOIJ1971 and the smallest influence of 2.1% is related to a nuclear reaction experiment 84Sr(d,p)85Sr MORTON1971.
III.2 King plot analysis
The isotope shift between two isotopes with mass numbers and , () is defined as the difference in transition frequencies: . The primary contributions to the isotope shift arise from the mass shift (MS) and field shift (FS). The MS results from the mass difference of isotopes and and is expressed as an electronic coefficient multiplied by the isotope-dependent reduced mass, given by: . The FS originates from the differing volumes of the two isotopes, factorized into an electronic, isotope-independent coefficient and the charge radius variance . The isotope shift composition in terms of electronic and nuclear quantities factorizes into the total isotope shift equation: , where the first term represents the MS, and the second term represents the FS king63; king1984isotope, with the dots denoting potential higher-order corrections and new physics contributions. The frequency shifts are commonly normalized by the reduced mass to obtain the modified isotope shift, . Consequently, the MS (the sum of the normal MS and specific MS) reduces to the electronic factor , while the FS factor is multiplied by the modified charge radius variance, , establishing a linear dependence between the two sets of modified frequency shifts known as the King linearity king63. To quantify the observed linearity, a measure of nonlinearity is defined flam18; Berengut18. A King plot analysis can be used to systematically quantify and visually examine isotope shifts in various atomic transitions referenced to the same isotope.
The refined mass-excess value obtained for 84Sr was employed to carry out an updated King plot analysis for two transitions at 589 nm and 598 nm, as detailed in miyake2019; Takano_2017. Figure 3 displays the King plot featuring both the literature mass values from AME2020 and our updated mass value. The main uncertainties in the plot are from the frequency shifts measurements in the transitions between 84Sr, 86Sr, and 87Sr relative to 88Sr as adopted from miyake2019. The transition of 84Sr-88Sr at 598 nm is from Takano_2017 which has a better precision than that of miyake2019. The observed nonlinearity from miyake2019, achieved through our new mass value for 84Sr is evident, and the uncertainty is no longer constrained by the mass uncertainty. Given that the uncertainty in the mass contributes to the calculation of the modified isotope shifts in similar ways, the impact on the linearity of the King plot is preserved due to the agreement of our result with the literature value.
A linear orthogonal distance regression analysis ODR2004 that provides unified standard error estimates for the uncertainties implanted in and directions, was conducted to determine the slope and intercept. The resulting slope and intercept from the fit to the AME2020 mass values and the updated mass value are 0.98148(51) and 8578(326) MHzamu, respectively. Our results align well with the fitting values of 0.981(5) and 8560(3450) MHzamu, reported in miyake2019, offering smaller uncertainties. These results confirm the nonlinearity announced in miyake2019, using the nonlinearity measure defined in flam18; Berengut18. The uncertainty from the mass values now allows for a potential future precision measurement of optical spectroscopy to reach a relative precision by more than three orders of magnitude, thus potentially leading to an unprecedented sensitivity for new physics. This will reach the spectroscopic precision necessary to attain state-of-the-art bounds on a spin-independent fifth force, as current calculations place the required uncertainty at < 1 Hz Ludlow15; Berengut18. This would essentially demand measuring optical atomic transitions with a relative precision of 10-15 or lower, a challenging accomplishment feasible only for a select few leading optical clock laboratories Ludlow15.
IV Conclusion and outlook
In summary, we have performed direct high-precision mass and double-beta decay value measurements of 84Sr using the PI-ICR technique with the JYFLTRAP double PTMS. The mass and values were determined with a precision approximately 30 times higher than those adopted in AME2020, aligning well with the literature. The new mass value was utilized to conduct an updated King plot analysis for transitions at 689 nm and 698 nm for 84,86,87Sr in relation to 88Sr. Our results showcase an enhanced nonlinear King plot, in good agreement with previously reported values, offering smaller uncertainties. The current contribution of the mass uncertainty to the King plot analysis enables future precision enhancements by several orders of magnitude in optical spectroscopy, mitigating statistical and systematic errors in both transitions. This potential advancement could lead to unparalleled sensitivity to new physics. More precise isotope shift measurements of Sr isotopes, providing stringent experimental constraints on King linearity, are highly required. Performing such a measurement with a radioactive isotope e.g., 90Sr, to avoid complications due to hyperfine structure, is feasible at the IGISOL facility. In conjunction with additional measurements of isotope shifts in elements such as calcium, this will contribute to enhancing precision in atomic structure calculations and imposing constraints on new physics.
Acknowledgements.
We acknowledge the staff of the Accelerator Laboratory of University of Jyväskylä (JYFL-ACCLAB) for providing stable online beam. We thank the support by the Academy of Finland under the Finnish Centre of Excellence Programme 2012-2017 (Nuclear and Accelerator Based Physics Research at JYFL) and projects No. 306980, No. 312544, No. 275389, No. 284516, No. 295207, No. 314733, No. 315179, No. 327629, No. 320062, No. 354589 and 345869. The support by the EU Horizon 2020 research and innovation program under grant No. 771036 (ERC CoG MAIDEN) is acknowledged. We express gratitude for the productive results stemming from conversations with R. de Groote and X. F. Yang.References
- King (1984) W. H. King, Isotope Shifts in Atomic Spectra (Plenum, New York, 1984).
- Flambaum et al. (2018) V. V. Flambaum, A. J. Geddes, and A. V. Viatkina, Phys. Rev. A 97, 032510 (2018).
- Dammalapati et al. (2009) U. Dammalapati, S. De, K. Jungmann, et al., Eur. Phys. J. D 53, 1 (2009).
- Shi et al. (2017) C. Shi, F. Gebert, C. Gorges, et al., Appl. Phys. B 123, 2 (2017).
- Nazé et al. (2015) C. Nazé, J. G. Li, and M. Godefroid, Phys. Rev. A 91, 032511 (2015).
- Frugiuele et al. (2017) C. Frugiuele, E. Fuchs, G. Perez, and M. Schlaffer, Phys. Rev. D 96, 015011 (2017).
- Berengut et al. (2018a) J. C. Berengut, D. Budker, C. Delaunay, V. V. Flambaum, C. Frugiuele, E. Fuchs, C. Grojean, R. Harnik, R. Ozeri, G. Perez, and Y. Soreq, Phys. Rev. Lett. 120, 091801 (2018a).
- Miyake et al. (2019) H. Miyake, N. C. Pisenti, P. K. Elgee, et al., Phys. Rev. Research 1, 033113 (2019).
- Manovitz et al. (2019) T. Manovitz, R. Shaniv, Y. Shapira, R. Ozeri, and N. Akerman, Phys. Rev. Lett. 123, 203001 (2019).
- Knollmann et al. (2019) F. W. Knollmann, A. N. Patel, and S. C. Doret, Phys. Rev. A 100, 022514 (2019).
- Counts et al. (2020a) I. Counts, J. Hur, D. P. L. Aude Craik, H. Jeon, C. Leung, J. C. Berengut, A. Geddes, A. Kawasaki, W. Jhe, and V. Vuletić, Phys. Rev. Lett. 125, 123002 (2020a).
- Solaro et al. (2020) C. Solaro, S. Meyer, K. Fisher, J. C. Berengut, E. Fuchs, and M. Drewsen, Phys. Rev. Lett. 125, 123003 (2020).
- Nesterenko et al. (2020) D. Nesterenko, R. de Groote, T. Eronen, Z. Ge, M. Hukkanen, A. Jokinen, and A. Kankainen, International Journal of Mass Spectrometry 458, 116435 (2020).
- Counts et al. (2020b) I. Counts, J. Hur, D. P. L. Aude Craik, H. Jeon, C. Leung, J. C. Berengut, A. Geddes, A. Kawasaki, W. Jhe, and V. Vuletić, Phys. Rev. Lett. 125, 123002 (2020b).
- Rehbehn et al. (2021) N.-H. Rehbehn, M. K. Rosner, H. Bekker, et al., Phys. Rev. A 103, L040801 (2021).
- Müller et al. (2021) R. A. Müller, V. A. Yerokhin, A. N. Artemyev, and A. Surzhykov, Phys. Rev. A 104, L020802 (2021).
- Hur et al. (2022) J. Hur, D. P. L. A. Craik, I. Counts, et al., Phys. Rev. Lett. 128, 163201 (2022).
- Figueroa et al. (2022) N. L. Figueroa, J. C. Berengut, V. A. Dzuba, et al., Phys. Rev. Lett. 128, 073001 (2022).
- Ono et al. (2022) K. Ono, Y. Saito, T. Ishiyama, et al., Phys. Rev. X 12, 021003 (2022).
- Stellmer et al. (2014) S. Stellmer, F. Schreck, and T. C. Killian, in Annual Review of Cold Atoms and Molecules, Vol. 2, edited by K. W. Madison, K. Bongs, L. D. Carr, A. M. Rey, and H. Zhai (World Scientific, Singapore, 2014) Chap. 1.
- Rana et al. (2012) R. Rana, M. Höcker, and E. G. Myers, Phys. Rev. A 86, 050502 (2012).
- Wang et al. (2021) M. Wang, W. Huang, F. Kondev, G. Audi, and S. Naimi, Chinese Physics C 45, 030003 (2021).
- Kondev et al. (2021) F. Kondev, M. Wang, W. Huang, S. Naimi, and G. Audi, Chinese Physics C 45, 030001 (2021).
- Eronen et al. (2012) T. Eronen, V. S. Kolhinen, V. V. Elomaa, D. Gorelov, U. Hager, J. Hakala, A. Jokinen, A. Kankainen, P. Karvonen, S. Kopecky, I. D. Moore, H. Penttilä, S. Rahaman, S. Rinta-Antila, J. Rissanen, A. Saastamoinen, J. Szerypo, C. Weber, and J. Äystö, European Physical Journal A 48, 1 (2012).
- Eronen et al. (2016) T. Eronen, A. Kankainen, and J. Äystö, Progress in Particle and Nuclear Physics 91, 259 (2016).
- Moore et al. (2013) I. D. Moore, T. Eronen, D. Gorelov, J. Hakala, A. Jokinen, A. Kankainen, V. S. Kolhinen, J. Koponen, H. Penttilä, I. Pohjalainen, M. Reponen, J. Rissanen, A. Saastamoinen, S. Rinta-Antila, V. Sonnenschein, and J. Äystö, Nuclear Instruments and Methods in Physics Research, Section B: Beam Interactions with Materials and Atoms 317, 208 (2013).
- Karvonen et al. (2008) P. Karvonen, I. D. Moore, T. Sonoda, T. Kessler, H. Penttilä, K. Peräjärvi, P. Ronkanen, and J. Äystö, Nuclear Instruments and Methods in Physics Research, Section B: Beam Interactions with Materials and Atoms 266, 4794 (2008).
- Nieminen et al. (2001) A. Nieminen, J. Huikari, A. Jokinen, J. Äystö, P. Campbell, and E. C. Cochrane, Nuclear Instruments and Methods in Physics Research, Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 469, 244 (2001).
- Savard et al. (1991) G. Savard, S. Becker, G. Bollen, H. J. Kluge, R. B. Moore, T. Otto, L. Schweikhard, H. Stolzenberg, and U. Wiess, Physics Letters A 158, 247 (1991).
- Nesterenko et al. (2018) D. A. Nesterenko, T. Eronen, A. Kankainen, L. Canete, A. Jokinen, I. D. Moore, H. Penttilä, S. Rinta-Antila, A. de Roubin, and M. Vilen, European Physical Journal A 54, 0 (2018).
- Nesterenko et al. (2021) D. A. Nesterenko, T. Eronen, Z. Ge, A. Kankainen, and M. Vilen, Eur. Phys. J. A 57, 302 (2021).
- Eliseev et al. (2014) S. Eliseev, K. Blaum, M. Block, A. Dörr, C. Droese, T. Eronen, M. Goncharov, M. Höcker, J. Ketter, E. M. Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, Applied Physics B: Lasers and Optics 114, 107 (2014).
- Eliseev et al. (2013) S. Eliseev, K. Blaum, M. Block, C. Droese, M. Goncharov, E. Minaya Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, Physical Review Letters 110, 82501 (2013).
- Kramida et al. (2020) A. Kramida, Yu. Ralchenko, J. Reader, and and NIST ASD Team, NIST Atomic Spectra Database (ver. 5.8), [Online]. Available: https://physics.nist.gov/asd [2021, January 19]. National Institute of Standards and Technology, Gaithersburg, MD. (2020).
- Kellerbauer et al. (2003) A. Kellerbauer, K. Blaum, G. Bollen, F. Herfurth, H. J. Kluge, M. Kuckein, E. Sauvan, C. Scheidenberger, and L. Schweikhard, European Physical Journal D 22, 53 (2003).
- Roux et al. (2013) C. Roux, K. Blaum, M. Block, C. Droese, S. Eliseev, M. Goncharov, F. Herfurth, E. M. Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, The European Physical Journal D 67, 1 (2013).
- Birge (1932) R. T. Birge, Physical Review 40, 207 (1932).
- Huang et al. (2021) W. Huang, M. Wang, F. Kondev, G. Audi, and S. Naimi, Chinese Physics C 45, 030002 (2021).
- Kellerbauer et al. (2007) A. Kellerbauer, G. Audi, D. Beck, K. Blaum, G. Bollen, C. Guénaut, F. Herfurth, A. Herlert, H.-J. Kluge, D. Lunney, S. Schwarz, L. Schweikhard, C. Weber, and C. Yazidjian, Phys. Rev. C 76, 045504 (2007).
- Haettner et al. (2011) E. Haettner, D. Ackermann, G. Audi, K. Blaum, M. Block, S. Eliseev, T. Fleckenstein, F. Herfurth, F. P. Heßberger, S. Hofmann, J. Ketelaer, J. Ketter, H.-J. Kluge, G. Marx, M. Mazzocco, Y. N. Novikov, W. R. Plaß, S. Rahaman, T. Rauscher, D. Rodríguez, H. Schatz, C. Scheidenberger, L. Schweikhard, B. Sun, P. G. Thirolf, G. Vorobjev, M. Wang, and C. Weber, Phys. Rev. Lett. 106, 122501 (2011).
- Booij et al. (1971) H. Booij, E. Van Hoek, H. Van Der Molen, W. Slot, and J. Blok, Nuclear Physics A 160, 337 (1971).
- Morton et al. (1971) J. Morton, W. Davies, W. Mclatchie, D. W., and J. Kitching, Nuclear Physics A 161, 228 (1971).
- Takano et al. (2017) T. Takano, R. Mizushima, and H. Katori, Applied Physics Express 10, 072801 (2017).
- King (1963) W. H. King, J. Opt. Soc. Am. 53, 638 (1963).
- Berengut et al. (2018b) J. C. Berengut, D. Budker, C. Delaunay, V. V. Flambaum, C. Frugiuele, E. Fuchs, C. Grojean, R. Harnik, R. Ozeri, G. Perez, and Y. Soreq, Phys. Rev. Lett. 120, 091801 (2018b).
- York et al. (2004) D. York, N. M. Evensen, M. L. Martınez, and J. De Basabe Delgado, American Journal of Physics 72, 367 (2004).
- Ludlow et al. (2015) A. D. Ludlow, M. M. Boyd, J. Ye, E. Peik, and P. O. Schmidt, Rev. Mod. Phys. 87, 637 (2015).